Test the null hypothesis Ho : u = 3.0against the alternative hypothesis HA: U < 3.0 , based on a random sample of 25 observat9
Answer #1

Solution :

\mu= 3.0

\bar x =2.8

\sigma =0.70

n = 25

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 3.0

Ha : \mu <3.0

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (2.8-3.0) / 0.70 / \sqrt 25

=−1.429

P(z < -1.43) = 0.0764

P-value = 0.077,

\alpha = 0.10

P-value < \alpha

p=0.077,<0.10, it is concluded that the null hypothesis is rejected.

when

\alpha = 0.05

P-value < \alpha

p=0.077,>0.05, it is concluded that the null hypothesis is not rejected.